The discriminant of -3x^2-5x=-4(correct to within .0001) is discriminant= _.
The number of real solutions of -3x^2-5x=-4 is_.

Answer by stanbon(75887)

The discriminant of -3x^2-5x=-4(correct to within .0001) is discriminant= _.
The number of real solutions of -3x^2-5x=-4 is_.
b^2-4ac = 5^2-4*-3*4=73
Since the discriminant is greater than zero there are
two unequal Real solutions.
========================